# LatentLens: Revealing Highly Interpretable Visual Tokens in LLMs

### v2

@@ Appendix D Ablations
Table 1: Ablation results showing LatentLens interpretability change and nearest-neighbor overlap with baseline, averaged across all layers. $\Delta$ Interp.: Change in % interpretable (baseline = 71.3%). Top-5 NN Overlap: Average number of matching neighbors out of 5. Token = same subword; Phrase = same full caption. Captioning ablations maintain interpretability with partial overlap ($\sim$2/5), suggesting multiple valid mappings exist.
@@
Baseline (OLMo + ViT-L) 71.3%
Different seed +1.3% 2.5 2.2
Linear connector +0.8% 2.1 2.0
First-sentence captions -1.6% 1.8 1.5
Unfrozen LLM +6.4% 1.9 1.5
Spatial Task (frozen) -33.2% 0.0 0.0
Spatial Task (unfrozen) -29.2% 0.0 0.0
@@
For each ablation we measure two metrics, averaged across all layers: (1) LatentLens Interpretability: The percentage of visual tokens judged interpretable by GPT-5, reported as change from baseline (71.3%). (2) LatentLens Overlap: Average number of matching top-5 LatentLens nearest neighbors with the baseline model (out of 5). We report both token-level overlap (same subword in both top-5 sets) and phrase-level overlap (same full contextual phrase).
@@
Results are summarized in Table˜1. Simply changing the seed results in only around half of the top-5 LatentLens results to overlap with the original run. All of them are still highly similar concepts but it is some indication that even the seed can have some influence, albeit it is unclear how semantically meaningful it is. On the other hand, we find encouraging results when replacing the 3-layer MLP connector with a single matrix (linear layer) and when replacing highly detailed captions as training data with single-sentence captions: The level of interpretability is almost the same, and we still see a significant amount of overlap between the original top-5 nearest neighbors (2.1 and 1.8 out of 5, respectively). Thus, the relationship between vision encoder and LLM embeddings spaces can be linearly aligned in an interpretable manner and moreover, short captioning data is enough to do so [^10] Follow-up work could explore what the limit of simplicity of the training data is for this interpretable alignment to appear.. Next, we observe a large increase in the amount of interpretable visual tokens ($6.4\%$) when unfreezing the LLM during training. Thus, with this more “expressive connector” in the form of some LLM weights, the model can learn a more interpretable alignment. Finally, we observe that the language-based task of captioning is necessary for interpretable visual tokens. When instead training the model to generate a single token (“top” or “bottom”) given a question about the location of an object, interpretability drops by around $30\%$, with zero overlap to the original LatentLens top-5 NNs. The tokens that are still marked as interpretable for this spatial task are the same few generic words such as “upper”, “background”, or “outside”.
@@
Appendix F Layer Alignment Details
@@
Figure˜11 shows the cosine similarity between each token at layer $\ell$ and its own representation at layer 0, averaged across all tokens. For text tokens, similarity to the input embedding drops rapidly within the first few layers (often below 0.4 by layer 4), reflecting the contextualization process where tokens incorporate information from surrounding context. In contrast, visual tokens maintain much higher similarity to their input representations—often above 0.8 through mid-layers—indicating they undergo substantially less transformation during LLM processing.
@@
Appendix G Results for an off-the-shelf model
@@
Figure˜12 shows the layer alignment and token drift for Qwen2-VL. We observe the expected diagonal pattern for most layers, where visual tokens at layer $n$ have most nearest neighbors from LLM layer $n$. However, for visual tokens at layer 0, we again observe the “leap” where most nearest neighbors come from LLM layer $4$. Likewise, visual tokens undergo less change throughout LLM layers compared to text tokens, similar to our controlled setup (Figure˜11), though they do change more in Qwen2-VL than in our frozen LLM models.
@@
Appendix H Fine-grained Interpretation Analysis
@@ H.1 Interpretation Types
Figure˜13 provides a visual breakdown of interpretation types aggregated across all models. Figure˜14 shows the breakdown for all 9 model combinations individually. We categorize LatentLens interpretations as:
@@
Figure˜15 shows the same analysis for Qwen2-VL-7B-Instruct, the off-the-shelf model from Appendix˜G. The pattern is similar: concrete interpretations dominate (62–78%), though we observe a slight decrease in later layers (26–27: 62%) with a corresponding increase in abstract interpretations. This may reflect Qwen2-VL’s unfrozen LLM having learned some layer-wise abstraction.
@@
H.2 Parts-of-Speech and Visual Attributes
Figure˜16 shows the distribution of parts-of-speech among all nearest neighbors for each of the 9 trained model combinations. Nouns dominate at approximately 45–50%, which aligns with the visual nature of the interpretations—visual tokens primarily encode objects and entities. Proper nouns account for 10–20%, verbs 10–15%, and adjectives around 5%. The distribution is relatively stable across layers, with some variation between models (e.g., OLMo-7B + ViT-L shows more variability). Figure˜17 shows the same analysis for Qwen2-VL-7B-Instruct, which exhibits similar patterns.
Figure˜18 shows the frequency of visual attribute words (colors, shapes, textures) for each trained model. Color words are the most common at around 5–6% in early layers, declining to around 3% in later layers. This suggests that raw color information is more prominent in early visual token representations. Shape and texture words are rare throughout (<<1%). Figure˜19 shows the same for Qwen2-VL.
@@ H.2 Parts-of-Speech and Visual Attributes
Appendix I Phrase-Level Interpretation Examples
@@
Appendix J Dynamic Corpus Generation
@@
We use an evolutionary search approach: given a visual token and its top-5 LatentLens nearest neighbors from the fixed corpus, we iteratively generate variations using GPT-5 and keep those with highest cosine similarity to the visual embedding. Crucially, since LLMs are autoregressive, we only modify words before the target token (the token whose contextual embedding we extract), keeping the target token at the end of each phrase.
@@ Results.
Appendix K Captioning Quality Evaluation
@@
Figure˜23 shows example captions from three representative models on the same validation image, illustrating how different model combinations produce captions of varying quality and detail.
@@
Appendix L Qualitative examples
@@
Appendix M Behind The Scenes
The goal of this section is to make science more transparent and engaging, showing not just the polished paper at the end but also all the detours and lessons learned. Note: Due to anonymity certain details might be omitted, and released at full length for the camera-ready version.
M.1 From start to finish

### v3

@@ Appendix D Ablations
Table 1: Ablation results showing LatentLens interpretability change and nearest-neighbor overlap with baseline, averaged across all layers. $\Delta$ Interp.: Change in % interpretable (baseline = 72.3%). Top-5 NN Overlap: Average number of matching neighbors out of 5. Token = same subword; Phrase = same full caption. Captioning ablations maintain interpretability with partial overlap ($\sim$2/5), suggesting multiple valid mappings exist.
@@
Baseline (OLMo + ViT-L) 72.3%
Different seed +0.4% 2.5 2.2
Linear connector -0.2% 2.1 2.0
First-sentence captions -2.5% 1.8 1.5
Unfrozen LLM +5.5% 1.9 1.5
Spatial Task (frozen) -34.2% 0.0 0.0
Spatial Task (unfrozen) -30.2% 0.0 0.0
@@
For each ablation we measure two metrics, averaged across all layers: (1) LatentLens Interpretability: The percentage of visual tokens judged interpretable by GPT-5, reported as change from baseline (72.3%). (2) LatentLens Overlap: Average number of matching top-5 LatentLens nearest neighbors with the baseline model (out of 5). We report both token-level overlap (same subword in both top-5 sets) and phrase-level overlap (same full contextual phrase).
@@
Results are summarized in Table˜1. Simply changing the seed results in only around half of the top-5 LatentLens results to overlap with the original run. All of them are still highly similar concepts but it is some indication that even the seed can have some influence, albeit it is unclear how semantically meaningful it is. On the other hand, we find encouraging results when replacing the 3-layer MLP connector with a single matrix (linear layer) and when replacing highly detailed captions as training data with single-sentence captions: The level of interpretability is almost the same, and we still see a significant amount of overlap between the original top-5 nearest neighbors (2.1 and 1.8 out of 5, respectively). Thus, the relationship between vision encoder and LLM embeddings spaces can be linearly aligned in an interpretable manner and moreover, short captioning data is enough to do so [^10] Follow-up work could explore what the limit of simplicity of the training data is for this interpretable alignment to appear.. Next, we observe a notable increase in the amount of interpretable visual tokens ($5.5\%$) when unfreezing the LLM during training. Thus, with this more “expressive connector” in the form of some LLM weights, the model can learn a more interpretable alignment. Finally, we observe that the language-based task of captioning is necessary for interpretable visual tokens. When instead training the model to generate a single token (“top” or “bottom”) given a question about the location of an object, interpretability drops by around $30\%$, with zero overlap to the original LatentLens top-5 NNs. The tokens that are still marked as interpretable for this spatial task are the same few generic words such as “upper”, “background”, or “outside”.
@@
Appendix F Cosine Similarity Distributions
Figure˜11 shows the distribution of top-1 nearest neighbor cosine similarities between visual tokens and their closest contextual text embeddings at layer 16 (a representative mid-layer). Several model combinations exhibit bimodal distributions with two distinct peaks. This pattern is most pronounced for LLaMA3-8B + CLIP and Qwen2-7B + CLIP, where one peak appears around 0.15–0.2 and a second around 0.35–0.4. A similar bimodal pattern is visible for Qwen2-7B + DINOv2.
In contrast models paired with SigLIP tend to show more unimodal distributions. The bimodal structure suggests that visual tokens may fall into two categories: those that align well with text representations (higher similarity) and those that remain more distant from the text embedding space (lower similarity). This could correspond to tokens representing salient visual content versus background or less semantically meaningful regions.
Figure 11: Distribution of top-1 nearest neighbor cosine similarities at layer 16. Each panel shows the histogram of cosine similarities between visual tokens and their closest contextual text embedding for one LLM–vision encoder combination. Several combinations (notably LLaMA3 + CLIP, Qwen2 + CLIP) exhibit bimodal distributions, suggesting distinct populations of visual tokens with different degrees of alignment to the text embedding space.
Appendix G Layer Alignment Details
@@
Figure˜12 shows the cosine similarity between each token at layer $\ell$ and its own representation at layer 0, averaged across all tokens. For text tokens, similarity to the input embedding drops rapidly within the first few layers (often below 0.4 by layer 4), reflecting the contextualization process where tokens incorporate information from surrounding context. In contrast, visual tokens maintain much higher similarity to their input representations—often above 0.8 through mid-layers—indicating they undergo substantially less transformation during LLM processing.
@@
Appendix H Results for an off-the-shelf model
@@
Figure˜13 shows the layer alignment and token drift for Qwen2-VL. We observe the expected diagonal pattern for most layers, where visual tokens at layer $n$ have most nearest neighbors from LLM layer $n$. However, for visual tokens at layer 0, we again observe the “leap” where most nearest neighbors come from LLM layer $4$. Likewise, visual tokens undergo less change throughout LLM layers compared to text tokens, similar to our controlled setup (Figure˜12), though they do change more in Qwen2-VL than in our frozen LLM models.
@@
Appendix I Fine-grained Interpretation Analysis
@@ I.1 Interpretation Types
Figure˜14 provides a visual breakdown of interpretation types aggregated across all models. Figure˜15 shows the breakdown for all 9 model combinations individually. We categorize LatentLens interpretations as:
@@
Figure˜16 shows the same analysis for Qwen2-VL-7B-Instruct, the off-the-shelf model from Appendix˜H. The pattern is similar: concrete interpretations dominate (62–78%), though we observe a slight decrease in later layers (26–27: 62%) with a corresponding increase in abstract interpretations. This may reflect Qwen2-VL’s unfrozen LLM having learned some layer-wise abstraction.
@@
I.2 Parts-of-Speech and Visual Attributes
Figure˜17 shows the distribution of parts-of-speech among all nearest neighbors for each of the 9 trained model combinations. Nouns dominate at approximately 45–50%, which aligns with the visual nature of the interpretations—visual tokens primarily encode objects and entities. Proper nouns account for 10–20%, verbs 10–15%, and adjectives around 5%. The distribution is relatively stable across layers, with some variation between models (e.g., OLMo-7B + ViT-L shows more variability). Figure˜18 shows the same analysis for Qwen2-VL-7B-Instruct, which exhibits similar patterns.
Figure˜19 shows the frequency of visual attribute words (colors, shapes, textures) for each trained model. Color words are the most common at around 5–6% in early layers, declining to around 3% in later layers. This suggests that raw color information is more prominent in early visual token representations. Shape and texture words are rare throughout (<<1%). Figure˜20 shows the same for Qwen2-VL.
@@ I.2 Parts-of-Speech and Visual Attributes
Appendix J Phrase-Level Interpretation Examples
@@
Appendix K Dynamic Corpus Generation
@@
We use an evolutionary search approach: given a visual token and its top-5 LatentLens nearest neighbors from the fixed corpus, we iteratively generate variations using GPT-4o and keep those with highest cosine similarity to the visual embedding. Crucially, since LLMs are autoregressive, we only modify words before the target token (the token whose contextual embedding we extract), keeping the target token at the end of each phrase.
@@ Results.
Appendix L Captioning Quality Evaluation
@@
Figure˜24 shows example captions from three representative models on the same validation image, illustrating how different model combinations produce captions of varying quality and detail.
@@
Appendix M Qualitative examples
@@
Appendix N Behind The Scenes
The goal of this section is to make science more transparent and engaging, showing not just the polished paper at the end but also all the detours and lessons learned.
N.1 From start to finish